OpenStudy (anonymous):

factor completely. ab-a-b+1

OpenStudy (jamesj):

ab - a - b + 1 = a (b - 1) - (b-1) = a(b - 1) + -1 (b - 1) = (a + -1)(b - 1) = (a - 1)(b - 1)

OpenStudy (anonymous):

thanks